If 2 years are added to 2/3 age of rani. She will be 3 years older then today.what is rani 's present age?
Answer by stanbon(75887)

If 2 years are added to 2/3 age of rani. She will be 3 years older then today.what is rani 's present age?
------
(2/3)r + 2 = r + 3
------
(1/3)r = -1
r = -3
----
That doesn't make any sense.
Please check your post.
Cheers,
Stan H.
-------